# Dynamic LLM-Agent Network: An LLM-agent Collaboration Framework with Agent Team Optimization

### v1

@@ Abstract
Large language model (LLM) agents have been shown effective on a wide range of tasks, and by ensembling multiple LLM agents, their performances could be further improved. Existing approaches employ a fixed set of agents to interact with each other in a static architecture, which limits their generalizability to various tasks and requires strong human prior in designing these agents. In this work, we propose to construct a strategic team of agents communicating in a dynamic interaction architecture based on the task query. Specifically, we build a framework named Dynamic LLM-Agent Network (DyLAN) for LLM-agent collaboration on complicated tasks like reasoning and code generation. DyLAN enables agents to interact for multiple rounds in a dynamic architecture with inference-time agent selection and an early-stopping mechanism to improve performance and efficiency. We further design an automatic agent team optimization algorithm based on an unsupervised metric termed Agent Importance Score, enabling the selection of best agents based on the contribution each agent makes. Empirically, we demonstrate that DyLAN performs well in both reasoning and code generation tasks with reasonable computational cost. DyLAN achieves 13.0% and 13.3% improvement on MATH and HumanEval, respectively, compared to a single execution on GPT-35-turbo. On specific subjects of MMLU, agent team optimization in DyLAN increases accuracy by up to 25.0%.  [^1] Code is available at https://github.com/SALT-NLP/DyLAN.
Figure 1: Overview of DyLAN. The top part shows how DyLAN outputs the answer in a feed-forward manner (Section 3). Inference-time agent selection is implemented with an LLM-empowered ranker (bottom-left), during which the low-performing agent is deactivated in subsequent time steps. In agent team optimization (Section 3.3), first the contribution of each agent is automatically evaluated using Agent Importance Score, denoted as ${\bm{I}}$. Then the top-ranked agents based on ${\bm{I}}$ will be selected as the optimized team of agents (bottom-right).
@@ 1 Introduction
Large language model (LLM) agents (Richards & et al., 2023; Nakajima, 2023; Reworkd, 2023) have achieved promising performance on a wide range of tasks, ranging from reasoning (Yao et al., 2023) to code generation (Shinn et al., 2023), and even to embodied tasks such as video gaming (Wang et al., 2023) and autopilot (Jin et al., 2023). Recent works tackle complicated tasks by combining different LLM agents in a team to collaborate on the same query (Du et al., 2023; Wang et al., 2023b; Jiang et al., 2023; Shinn et al., 2023; Zheng et al., 2023; Wu et al., 2023). In these works, such LLM-agent team is often set up statically with predefined roles for each agent, such as introducing programmer and tester to communicate sequentially in a fixed order to improve code generation (Dong et al., 2023).
In the case of collaborative work between different LLM agents on a specific task, a static setup with predefined roles can, however, have several problems: (1) Manual setups require us to design task-specific roles for LLM agents in each specific domain (Dong et al., 2023), making it difficult to generalize to other domains or distinct tasks. (2) Additionally, under a static setup, LLM agents interact to produce answers in a fixed order (Shinn et al., 2023) or in one single round (Jiang et al., 2023), which might result in sensitivity. (3) Predefined roles require strong human priors to design, and may not well align with the actual situations. In addition, there is no systematic way to ensure that the LLM-agent collaboration systems have a sufficient number of LLM agents and, most importantly, an optimized team of agents.
The landscape of LLM-agent collaborations requires a systematic framework in order to improve generalizability, efficiency, and performance. Therefore, several properties should be exhibited: (1) Task Agnostic: Prior works have suffered from the difficulty of generalizing across various domains due to the dependence on task-specific tools. A task-agnostic system can facilitate the fast adaptation of existing approaches to new situations. (2) Efficient: Instead of assigning agents in a static pattern, dynamically removing agents with uninformative responses can prevent the creation of useless information as well as ensure accuracy in the process of reaching consensus. (3) Agent Team Optimization: With thousands of open-source and unlimited LLM-generated prompts that serve a variety of roles, it is difficult to identify what the optimal team of agents might be. Ideally, multi-agent systems should be able to adapt their composition in response to the particular domain of a query with minimal supervision. While a few efforts have been made (Chen et al., 2023b; Ruan et al., 2023; Zhang et al., 2023; Wang et al., 2023b; Lu et al., 2023; Aggarwal et al., 2023; Liu et al., 2023b; Besta et al., 2023), a framework that integrates all of these aspects is still lacking.
To fill these gaps, in this work, we introduce a framework named Dynamic LLM-Agent Network (DyLAN) in Section 3. Concretely, we employ a feed-forward network to formulate the process of LLM-agent collaboration for arbitrary tasks. As a result, we view LLM agents at specific time steps as nodes in a network and the messages they exchange at different time steps as edges. In this way, we can organize multi-round LLM-agent collaboration into a multilayered network. To enhance the efficiency and the performance of LLM-agent collaboration on various tasks, we use an LLM-empowered ranker (Qin et al., 2023) to rank different LLM agents and deactivate low-performing agents in the subsequent interaction (i.e., inference-time agent selection), thereby creating a dynamic architecture of interactions. Additionally, we introduce an early-stopping mechanism via a Byzantine Consensus to ensure efficiency, by terminating the inference process when agents within the same layer reach consensus. Moreover, we propose automatic agent team optimization without strong human priors to increase the efficiency and effectiveness of agent team initialization. Specifically, to select the top-$k$ agents, we introduce a three-step procedure where we first ask each agent to rate its predecessors on their solutions (propagation), and then for each agent, aggregate the ratings from their successors to quantify its contribution (aggregation). Finally, after summing up the ratings across all time steps, we derive an Agent Importance Score for each agent. Then, we can select the top-performing agents based on their importance scores to obtain the optimized team of agents (selection). In this way, DyLAN achieves the task-agnostic property by formulating LLM-agent collaboration into a feed-forward network to decouple the interaction architecture and task-specific design, exhibits efficiency through inference-time agent selection and early-stopping mechanism, and enables agent team optimization through a selection algorithm based on agent importance scores. We evaluate DyLAN on multiple representative tasks such as general reasoning, arithmetic reasoning, and code generation, and find that DyLAN demonstrates higher accuracy and efficiency over baselines. We also show that Agent Importance Score can be used as a solid unsupervised indicator for optimizing the composition of agents for DyLAN. On specific subjects of MMLU, agent team optimization in DyLAN increases accuracy by up to 25.0%.
To sum up, our contributions are as follows:
  • We propose a generic LLM-agent collaboration framework DyLAN that organizes agents into a multi-layered feed-forward network with dynamic architecture by introducing inference-time agent selection and early-stopping mechanism.
  • We develop an automatic agent team optimization algorithm for DyLAN based on an unsupervised Agent Importance Score, in order to achieve both efficiency and effectiveness.
  • Empirically, DyLAN demonstrates high accuracy, efficiency, and stability in general reasoning, arithmetic reasoning, and code generation tasks.
@@ 2 Related Work
Interaction Architecture in LLM-agent Collaboration
Collaboration between multiple LLM agents has demonstrated strong performance on a variety of tasks in recent years and has emerged as a promising approach to enhance the capabilities of individual LLMs. To enable collaborations between multiple agents, recent studies have developed different interaction architectures and assigned agents in static patterns. For instance, Du et al. (2023); Liang et al. (2023); Xiong et al. (2023) take multiple LLM instances into a debate for a fixed number of rounds, thus boosting their factuality and reasoning capacities. Instead of calling LLMs iteratively, Ning et al. (2023) distribute LLMs in parallel and concatenates their answers to produce better results. To aggregate multiple LLM responses, Jiang et al. (2023) generate candidates by different LLMs in one round and uses pairwise ranking to combine the top responses into one final response. It is worth noting that Hao et al. (2023) organizes LLM instances into linear layers and adopts supervised learning in context space, not the scenario in which we are interested. However, running LLMs in a static architecture may limit their performance and generalization. Although Zhang et al. (2023) adopt a dynamic directed acyclic graph structure during inference, they merely focus on reasoning and are incompatible with external tools and diverse agents with different roles. In this work, we propose an interaction architecture that can be adjusted dynamically based on the query and be compatible with feedback and tool augmentation.
Evaluation of the Contribution of LLM Agents
It is non-trivial to evaluate the contribution of each LLM agent in a multi-agent system, especially when they communicate over multiple rounds. In the single-round setting, given multiple candidates to select the best one, existing methods use LLMs heavily for evaluation. However, Xiong et al. (2023) show that LLMs tend to be overconfident. For more reliable results, pairwise ranking based on an additional LLM-powered ranker has been introduced in Jiang et al. (2023) for greater accuracy. To rank $n$ responses with an independent LLM in a single round, they compare all $O(n^{2})$ pairs. For better efficiency, Qin et al. (2023) use a $k$-length sliding window to choose top $k$ responses within $O(nk)$ pairwise comparisons. However, these methods have not been extended to multi-round settings. Inspired by the neuron importance score (Yu et al., 2018), we evaluate agents by propagating and aggregating single-round peer ratings. In this way, we then introduce an unsupervised metric called Agent Importance Score to quantify the contribution of each agent in multi-round collaborations (Section 3.3).
Team Optimization for LLM Agents
@@
In terms of designing and selecting agents, Ruan et al. (2023) decompose tasks to choose or create tools accordingly. Using LLM as a planner, Lu et al. (2023) sequentially select agents and tools according to their descriptions. Wang et al. (2023b) use LLMs to generate prompts for agents in response to a task query. During inference, Chen et al. (2023b) select a fixed number of agents from a set of manual prompt candidates via an additional LLM during each round of discussion. However, manual prompts require careful design, and predefined or generated descriptions may not result in the desired abilities of the agents during inference. Therefore, selecting an appropriate team of agents based on their response to a query is necessary and more appropriate. While team optimization for LLM agents is a relatively new area, human-team optimization has been studied for a long time. For instance, Liu et al. (2015) show that skill contribution is essential for selecting crowd workers to solve outsourced tasks efficiently. Based on peer rating, Lykourentzou et al. (2022) develop an algorithm for managing online workers in an optimal organization. Building upon these prior works, we introduce an automatic algorithm to optimize the team of agents by quantifying agents’ contributions based on their peer ratings.
Table 1: Comparison between DyLAN and representative previous works. In the second row, nodes denote agents at different time steps (${\mathbb{A}}$), arrows represent edges ($\mathbb{E}$), and color indicates the role of an agent. Among these works, DyLAN is the only one which demonstrates all four key dimensions of LLM-agent collaboration, i.e., compatible with multiple roles, having an early-stopping mechanism, supporting dynamic interactions, and performing agent team optimization.
@@
Method Single LLM-Blender PHP Reflexion LLM Debate DyLAN
Interaction Architecture
$({\mathbb{A}},\mathbb{E})$
Multiple Roles $\times$ $\times$ $\times$ $\times$
Early Stopping $\times$ $\times$ $\times$ $\times$
Dynamic Interactions $\times$ $\times$ $\times$ $\times$
Agent Team Optimization $\times$ $\times$ $\times$ $\times$ $\times$
3 Dynamic LLM-Agent Network (DyLAN)
We introduce an LLM-agent collaboration framework named Dynamic LLM-Agent Network (DyLAN) to allow dynamic interaction between agents in LLM-agent collaborations, as illustrated in Figure 1. We formulate the systems in feed-forward networks to represent interaction architecture (Section 3.1). Under the formulation, we elaborate DyLAN’s dynamic architecture designs in Section 3.2. Finally, we introduce agent team optimization, based on our unsupervised metric Agent Importance Score on top of DyLAN (Section 3.3).
3.1 Formulation
In LLM-agent collaborations, agents exchange textual messages in multi-round interactions. By viewing LLM agents at different time steps as nodes, we can treat the message exchange between multiple nodes as edges in a feed-forward network. Agents may send their reasoning steps to each other on reasoning tasks, or send code completions on code generation tasks, as illustrated in Appendix D.6. We then demonstrate how both the inference process and agent team optimization can be viewed as message passing algorithms over this abstract network. This formulation serves as an essential task-agnostic basis for the construction of DyLAN. According to the formulation, an LLM-agent collaboration comprises the following three components (let the input query be $q$):
Node The node denotes an agent at a specific time step. The agent takes the contexts from other agents at the previous time step as input and generates responses based on the input query. An agent can be (I) an LLM agent that could be augmented with tools, (II) an independent tool like a script or a dedicated model, or (III) a cluster of tools. Based on the definition, we focus on textual information exchange between agents. Formally, the $i$-th agent at the $t$-th time step can be represented as a function $a_{t,i}$ mapping the agent’s prompt $p_{i}$ (empty for (II) and (III) agents), the input query $q$, and responses from predecessor agents ${\mathbb{R}}_{t-1}$ to the response of itself: $r_{t,i}=a_{t,i}(p_{i},q,{\mathbb{R}}_{t-1})$, where ${\mathbb{R}}_{t-1}=\{r_{t-1,j}|j=1,2,...\}$. Let ${\mathbb{A}}$ be the set of all nodes, $n$ be the total amount of agents, and $T$ be the maximum time step.
Edge Edges refer to the communication channels between nodes in ${\mathbb{A}}$ during the multi-agent collaboration. Let $\mathbb{E}$ represents the set of all edges in the system. The edge is directional and could be formally represented as $e=(a_{t-1,i},a_{t,j})\in\mathbb{E}$, where $a_{t-1,i},a_{t,j}$ represents adjacent agents that could pass textual information: $a_{t,j}$ could perceive $a_{t-1,i}$’s output as its context. Thus, nodes are linked by edges to form the interaction architecture of agent collaboration systems, in shape of a feed-forward network $S=({\mathbb{A}},\mathbb{E})$.
Message Passing Message passing algorithms can guide information flow through the nodes and edges in the feed-forward network. In our LLM-agent based feed-forward network formulation, we envision two types of message passing. In a forward passing fashion, by passing messages to specific agents across different time steps, LLM-agent collaboration systems generate the final answer in response to the task query. Formally, we denote the output as $o$, which is the result of the query $q$ processed by system $S$ and an algorithm $f_{\mathrm{Infer}}$ denoting the inference process explicitly (Algorithm 1 for DyLAN): $o=f_{\mathrm{Infer}}(S,q)$. In a backward passing manner, The calculation process of Agent Importance Score ${\bm{I}}$ is calculated by an algorithm $f_{\mathrm{Imp}}$ to pass scores backward along edges (Algorithm 2 for DyLAN): ${\bm{I}}=f_{\mathrm{Imp}}(S,q)$.
The formulation decouples the interaction architecture and the algorithm of inference process. Using this formulation, we can categorize prior works based on their interaction architectures in Table 1.
3.2 Construction of DyLAN
With the aforementioned formulation, we step further to stack multiple layers along the temporal axis (e.g., different rounds of interaction) to build Dynamic LLM-Agent Network (DyLAN). To improve the efficiency and the performance of LLM-agent collaboration on various tasks, we propose to (1) add inference-time agent selection at a middle time step and (2) use Byzantine Consensus to terminate inference at a proper layer.
We define a layer as a set of agents functioning at the same time step along the temporal axis. In each layer, each node receives responses all other nodes from the previous layer, i.e., the previous time step. Such communications form edges between layers, as shown in Figure 1. The query is fed into nodes in the first layer. Therefore, when LLM agents diverge on a query, they receive others’ responses and have a debate by commenting on each other’s solutions, thus deepening the reasoning procedure and reaching better consistency at a certain time step with the vantage of multiple expertise. Agents assigned diverse roles in DyLAN might respond from their specific perspectives.
To focus on the best solutions and reduce the computational cost, we set up the inference-time agent selection at $L$-th layer by selecting the top-$m$ responses to feed forward. We use an additional LLM agent, as the “LLM Ranker” in Figure 1, to analyze responses from the former layer and identify the best ones, following Jiang et al. (2023). Agents identified as not being useful are deactivated in subsequent layers, so as are the edges connecting them.

### v2

# A Dynamic LLM-Powered Agent Network for Task-Oriented Agent Collaboration

@@ Abstract
Recent studies show that collaborating multiple large language model (LLM) powered agents is a promising way for task solving. However, current approaches are constrained by using a fixed number of agents and static communication structures. In this work, we propose automatically selecting a team of agents from candidates to collaborate in a dynamic communication structure toward different tasks and domains. Specifically, we build a framework named Dynamic LLM-Powered Agent Network (DyLAN) for LLM-powered agent collaboration, operating a two-stage paradigm: (1) Team Optimization and (2) Task Solving. During the first stage, we utilize an agent selection algorithm, based on an unsupervised metric called Agent Importance Score, enabling the selection of best agents according to their contributions in a preliminary trial, oriented to the given task. Then, in the second stage, the selected agents collaborate dynamically according to the query. Empirically, we demonstrate that DyLAN outperforms strong baselines in code generation, decision-making, general reasoning, and arithmetic reasoning tasks with moderate computational cost. On specific subjects in MMLU, selecting a team of agents in the team optimization stage improves accuracy by up to 25.0% in DyLAN. [^1] Code is available at https://github.com/SALT-NLP/DyLAN.
Figure 1: DyLAN adopts a two-stage paradigm. Agents communicate in a structure of the temporal feed-forward network (T-FFN). At the “Team Optimization” stage, DyLAN performs agent selection for the most contributory agents in a primary collaboration, oriented to tasks or domains. The selected agents then collaborate dynamically for an answer on the given query at the “Task Solving” stage.
@@ 1 Introduction
Large Language Model (LLM) agents (Richards & et al., 2023; Nakajima, 2023; Reworkd, 2023) have demonstrated promising performance on various tasks, ranging from reasoning (Yao et al., 2023), code generation (Shinn et al., 2023) to embodied tasks such as video gaming (Wang et al., 2023a) and autopilot systems (Jin et al., 2023). Given the impracticality of a single agent managing all these tasks efficiently, recent research has shifted towards multi-agent collaborations, yielding significant advancements (Li et al., 2023; Du et al., 2023; Wang et al., 2023c; Jiang et al., 2023; Shinn et al., 2023; Chen et al., 2024; Wu et al., 2023).
As an analogy to human society, how human teams function may provide valuable insights for developing more effective multi-agent collaboration systems. For instance, recent studies have demonstrated that certain effective communication structures, derived from human society, also play a positive role in multi-agent collaborations (Yin et al., 2023; Chen et al., 2024). In addition to communication structures, another notable characteristic of human teams is that they would optimize team members according to the given task. Take medical consultations as an example. Collaborations in dynamic structures are evident when the composition of the team changes within the procedure of consultation, as some doctors may become less relevant in major as the conversation goes deeper and “leave” the consultation, leading to corresponding changes in the communication structure. Team optimization is often observed in the varying initial composition of doctors for consultations with different diseases, influenced by changes in the related medical fields and the contribution of each doctor. These characteristics prompt an important question: Does a dynamically changing team of agents benefit LLM-powered agent collaborations similarly?
However, the question is not well-addressed yet. While various communication structures have been studied for different tasks, such as debating for reasoning (Du et al., 2023; Liang et al., 2023; Xiong et al., 2023) and self-collaboration for coding (Dong et al., 2023; Qian et al., 2023a; b), these communication structures do not alter members in the agent team and remain fixed throughout the collaboration. It indicates that task-oriented dynamic selections in agents are not thoroughly explored in current research. Furthermore, in the context of agent teams, most existing studies opt for hand-crafting agents from human priors (Liu et al., 2023; Nakajima, 2023; Hong et al., 2024; Shinn et al., 2023; Li et al., 2023) or employ an LLM to generate them (Wang et al., 2023c; Chen et al., 2023b; Christianos et al., 2023). These approaches generally predefine agents without further validation of the collaboration process. This leads to static agent teams or rebuilding teams without principled verification (Chen et al., 2024). Challenges still remain for optimization methods.
As a first attempt towards addressing the above question, we introduce a novel framework named Dynamic LLM-Powered Agent Network (DyLAN). DyLAN conceptualizes multi-agent collaboration using temporal feed-forward networks (T-FFNs). In this formulation, each communication step of the agents corresponds to a network layer, with nodes representing the agents involved at that step and edges indicating communications between agents, for incorporating dynamic agent teams agnostically. DyLAN functions in two stages to incorporate task-oriented agent collaborations (Figure 1). The first stage is termed Team Optimization, where we select top contributory agents unsupervisedly among the initial team of candidates according to the task query, based on their individual contributions. We propose a forward-backward message passing algorithm on the T-FFN termed agent selection in Section 3.4, inspired by the back-propagation algorithm (Rumelhart et al., 1986) and neuron importance scores (Yu et al., 2018). This algorithm measures the contribution of each agent at the first stage with an unsupervised metric named Agent Importance Score. The most contributory agents form a smaller team to collaborate at the second stage — Task Solving, thereby minimizing the impact of less effective agents on the final answer. Specifically, the collaboration begins with a team of agents, and an LLM-powered ranker in the middle dynamically deactivates low-performing agents (i.e., agent team reformation), thus expanding the T-FFN, integrating dynamic communication structures into DyLAN (Section 3.3.2). Incorporating agent selection, DyLAN effectively identifies and coordinates a task-oriented team of agents in a principled way. Extensive experiments demonstrate that DyLAN outperforms strong baselines in various tasks, including code generation, decision-making, general reasoning, and arithmetic reasoning. Notably, agent selection in DyLAN has improved accuracy by up to 25.0% in certain subjects of the MMLU dataset (Hendrycks et al., 2021a), underlining the significance of dynamic agent teams.
In summary, our contributions are threefold:
  • We introduce a novel framework named DyLAN for task-oriented agent collaboration in two stages with agent selection, marking a significant advancement in the study of dynamic agent teams.
  • DyLAN innovatively formulates agent collaborations in temporal feed-forward networks with agent team reformation, enhancing its adaptability and reducing dependence on human preconceptions.
  • Empirical results demonstrate the superior accuracy, efficiency, and stability of DyLAN across various tasks, underscoring the need for dynamic agent teams.
@@ 2 Related Work
Team Optimization of LLM-Powered Agents The construction of agent teams is the essential and initial step for LLM-powered agent collaboration. TPTU (Ruan et al., 2023) and Chameleon (Lu et al., 2023) decompose tasks to choose or create tools accordingly. Recent studies also use LLMs to generate a fixed number of role prompts for agents in response to a task query (Wang et al., 2023c; Suzgun & Tauman Kalai, 2024), or for each round of discussion (Chen et al., 2024). However, manual prompts require careful design, which is impractical for adaptation on each task or domain, and prompting LLMs with predefined or generated descriptions may not result in the desired abilities of the agents without verification. Therefore, posteriorly selecting a team of agents based on their actual behaviors in the collaboration according to the task becomes essential. While team optimization for LLM agents is a relatively new area, human-team optimization has been studied for a long time. For instance, Liu et al. (2015) show that skill contribution is essential for selecting crowd workers to solve outsourced tasks efficiently. Based on peer rating, researchers have developed an algorithm for managing online workers in an optimal organization (Lykourentzou et al., 2022). Drawing inspirations, we introduce an unsupervised algorithm to select a team of agents by quantifying their contributions based on peer ratings in Section 3.4.
Communication Structures in LLM-Powered Agent Collaboration Collaboration between multiple LLM agents has demonstrated strong performance on various tasks in recent years and has emerged as a promising approach to enhance the capabilities of individual LLMs. To enable collaborations between multiple agents, recent studies have developed different communication structures and assigned agents in pre-defined architecture. For instance, researchers have found taking multiple LLM instances to debate for a fixed number of rounds can boost their factuality and reasoning capacities (Du et al., 2023; Liang et al., 2023; Xiong et al., 2023). To aggregate multiple LLM responses, LLM-Blender (Jiang et al., 2023) calls different LLMs in one round and uses pairwise ranking to combine the top responses. It has also been shown effective in distributing workloads to LLMs and concatenating their answers, thus producing better results (Ning et al., 2024; Suzgun & Tauman Kalai, 2024; Qiao et al., 2024). It is worth noting that existing studies (Hao et al., 2023; Zhang et al., 2023b) have tried organizing LLM instances into linear layers, but they mainly studied supervised learning in context space and LLM evaluation, respectively, not the scenario in which we are interested. However, running LLMs in a static architecture may limit the performance and generalization. On specific reasoning tasks, adopting a dynamic directed acyclic graph structure for LLMs has been shown effective (Zhang et al., 2023).Also, recent studies (Yin et al., 2023; Chen et al., 2024; Zhang et al., 2023a; Zhuge et al., 2024) have demonstrated that optimal communication structures vary with tasks and compositions of agents. Aligned with the findings, we propose a structure that adjusts dynamically based on selecting agents according to the tasks and the construction of the agent team in Section 3.3.2.
Evaluation of the Contribution of LLM-Powered Agents It is non-trivial to evaluate the contribution of each LLM agent in a multi-agent system, especially when they communicate over multiple rounds. In the single-round setting, existing methods use LLMs heavily for evaluation. To overcome the over confidence of LLMs (Xiong et al., 2024), pairwise ranking based on an additional LLM ranker has been introduced in LLM-Blender (Jiang et al., 2023). To rank $n$ responses with an independent LLM in a single round, they compare all $O(n^{2})$ pairs. For better efficiency, researchers use a $k$-length sliding window to choose top $k$ responses within $O(nk)$ pairwise comparisons (Qin et al., 2023). However, these methods have not been extended to multi-round settings. Inspired by the neuron importance score (Yu et al., 2018), we evaluate agents by propagating and aggregating single-round peer ratings in a back-propagation manner (Rumelhart et al., 1986). In this way, we then introduce an unsupervised metric called Agent Importance Score to quantify the contribution of each agent in multi-round collaborations (Section 3.4).
@@
Method Single Exec. LLM-Blender LLM Debate Reflexion CAMEL AgentVerse DyLAN
Communication Structure
$(\mathcal{V};E)$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Multiple Roles $\times$ $\times$ $\times$ Manual Manual Generated Man.&Gen.
Early Stopping $\times$ $\times$ $\times$
Dynamic Structure $\times$ $\times$ $\times$ $\times$ $\times$
Team Optimization $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Table 1: Comparison between DyLAN and representative previous works. In the second row, nodes denote agents at different time steps ($\mathcal{V}$), arrows represent edges ($E$), and color indicates the role of agents.
3 Dynamic LLM-Powered Agent Network
3.1 Overview
We introduce a framework for LLM-powered agent collaboration named Dynamic LLM-Powered Agent Network (DyLAN), facilitating dynamic communication structures and automatically task-oriented agent selection in a two-stage fashion (Figure 1): an optimized agent team is constructed in the first stage “Team Optimization” through a preliminary trial and then the team collaborates to solve the task in the second stage “Task Solving”.
A core component of DyLAN is the temporal feed-forward networks (T-FFNs), whose nodes denote agents and edges denote the communication channels between agents (Figure 2 left). T-FFNs serve not only as the abstraction of communication structures but also the computation graph. From this perspective, as shown in Table 1, various LLM-powered agent collaboration systems (Jiang et al., 2023; Shinn et al., 2023; Du et al., 2023; Li et al., 2023; Chen et al., 2024) can be represented by similar network structures as T-FFNs. DyLAN is the only framework that supports multiple agents with roles and tools, early stopping (Section 3.3.2), dynamic communication structures and team optimization simultaneously. To be specific, for team optimization, our agent selection algorithm is performed as a backward message passing algorithm on the T-FFN (Figure 2 right), and for task solving, agent team reformation expands the T-FFN dynamically with messages passing forward.
To make it easier to understand, we will start by explaining the formulation of T-FFNs. Then, we will move on to the task solving stage, and finally, we will explain the team optimization stage, which relies on components of task solving.
3.2 Temporal Feed-Forward Networks (T-FFNs)
Figure 2: The left part shows how DyLAN outputs the answer in a temporal feed-forward network (T-FFN), where nodes represent agents at specific time steps (Section 3.2). Agent team reformation functions in the middle steps, during which the low-performing agent is deactivated in subsequent time steps. The right part depicts agent selection (Section 3.4), where the contribution of each agent in a primary trial is automatically evaluated in three steps using Agent Importance Score, denoted as $\bm{I}$. Then, the top-ranked agents based on $\bm{I}$ will be selected as the optimized, task-oriented team of agents.
A T-FFN is a multi-layer network, of which each layer represents a time step. Its formal definition is as follows.
Definition 1 (Agents)
Agents participating the collaboration are represented by
$\mathcal{A}=\{a_{1},a_{2},\cdots,a_{N}\},$ (1)
where $N$ denotes the total number of agents, and $a_{i}$ can be (I) an LLM-powered agent possibly equipped with tools, or (II) an independent tool, e.g., scripts, code interpreters.
Definition 2 (Nodes)
The $t$-th layer of a T-FFN consists of $N$ nodes, each of which corresponding to one agent:
$\mathcal{V}_{t}=\{v_{t,1},\cdots,v_{t,N}\},$ (2)
where $t=1,\cdots,T$, and node $v_{t,i}$ corresponds to agent $a_{i}$.
Definition 3 (Edges)
Edges in a T-FFN refer to the communication channels between nodes, forming the communication structure between agents. Specifically, the set of edges between the nodes in layer $t-1$ and $t$ is denoted as
$E_{t-1,t}=\{(v_{t-1,i},v_{t,j})\}\subseteq\mathcal{V}_{t-1}\times\mathcal{V}_{t},$ (3)
where $t=2,\cdots,T$, and $(v_{t-1,i},v_{t,j})$ denotes an edge connecting nodes $v_{t-1,i}$ and $v_{t,j}$.
Definition 4 (T-FFN)
Finally, the T-FFN corresponding to the collaboration is defined as a $T$-layer network:
$\mathcal{G}=(\mathcal{V}_{1},\cdots,\mathcal{V}_{T};E_{1,2},\cdots,E_{T-1,T}).$ (4)
Note that we only consider T-FFNs where edges only exist in adjacent layers. However, edge can be added to arbitrary pairs of nodes, making the T-FFN capable of representing more complex communication structures.
3.3 Task Solving
Task solving involves performing inference on the T-FFNs, jointly with agent team reformation, which be elaborated on in the subsequent two sections.
3.3.1 Inference
Before details, we first introduce the formulation of message passing on T-FFNs.
Definition 5 (Message Passing)
Given a T-FFN, a node $v_{t,j}$, a set of adjacent nodes $\mathcal{U}=\{u_{1},\cdots,u_{K}\}$, and the messages =$\mathcal{M}=\{m_{u_{1}},\cdots,m_{u_{K}}\}$ received by $v_{t,j}$, where $K$ is the size of $\mathcal{U}$, and $m_{u_{k}}$ is the message sent from $u_{k}$ to $v_{t,j}$, message passing aggregates all the messages $\mathcal{M}$ to produce a updated message $\hat{m}_{v_{t,j}}$ for $v_{t,j}$, which is formally defined as
$\hat{m}_{v_{t,j}}=f_{\mathrm{mp}}\left(\mathcal{M},v_{t,j}\right),$ (5)
where $f_{\mathrm{mp}}(\cdot,\cdot)$ is the aggregator function.
Definition 6
We refer the algorithm as forward message passing when $\mathcal{U}$ is the set of all adjacent nodes of $v_{t,j}$ from the previous time step, i.e., $\mathcal{U}=\{u_{k}|\forall u_{k},(u_{k},v_{t,j})\in E_{t-1,t}\}$. Similarly, it is referred as backward message passing when $\mathcal{U}$ is the set of all adjacent nodes from the next time step: $\mathcal{U}=\{u_{k}|\forall u_{k},(v_{t,j},u_{k})\in E_{t,t+1}\}$.
With the above formulation, we can describe the inference process of a T-FFN $\mathcal{G}$ in the manner of forward message passing. During collaborations on a given task, an agent at a specific time step takes the responses, i.e., messages, from other agents at the previous time step as input and generates responses based on the task query. Based on different types of agents at $v_{t,j}$, we can implement $f_{\mathrm{mp}}(\cdot,v_{t,j})$ respectively: (I) concatenating input messages along with the task query into prompt templates (refer to task instructions in Appendix D) and take the response from LLM after generation or tool calling, or (II) filter the input that the tool can process, e.g., code completions and structured text.
During inference, we begin feeding the task query $q\in\mathcal{Q}$ into agents at time step 1 ($\mathcal{V}_{1}$), where $\mathcal{Q}$ denotes the dataset. By passing responses of nodes $\mathcal{V}_{t-1}$ at time step $t-1$ to nodes $\mathcal{V}_{t}$ at $t$, agents can perceive responses from all nodes at the previous time step and perform collaborative behavior, which might include criticizes, advice, refinement, or quality reviews, depending on the implementation of agents. Formally, the inference process is defined as
$f_{\mathrm{Infer}}(q,\mathcal{G})=o,$ (6)
where $o=\mathrm{argmax}\{\mathcal{M}_{T}\}$ and $\mathcal{M}_{T}$ denotes the responses from $\mathcal{V}_{T}$. Please refer to Algorithm 1 for detailed procedure.
3.3.2 Agent Team Reformation
Given a set of agents $\mathcal{A}$, agent team reformation aims to identify more contributory agents and construct a dynamic communication structure accordingly. To this end, we leverage an additional LLM instance, referred as the “LLM Ranker”, to analyze responses from the agents participate in the current time step and give out a ranking, prompted by the template of “Ranker” in Appendix D. Then, the top-ranked agents are allowed to participate in the next time step. In other words, edges will only be added for these top-ranked agents, resulting in a dynamic communication structure.
Formally, suppose the set of agents participates in time step $t$ is $\mathcal{A}_{t}=\{a_{k}\}$, and the top-ranked agents are $\mathcal{A}_{t+1}=\{a_{l}\}$, where $k$ and $l$ are the indices of the agents as defined in Equation 1, then we can obtain two nodes sets $\mathcal{V}_{t}^{\prime}=\{v_{t,k}|\forall a_{k}\in\mathcal{A}_{t}\}$ and $\mathcal{V}_{t+1}^{\prime}=\{v_{t+1,l}|\forall a_{l}\in\mathcal{A}_{t+1}\}$, and the edge set $E_{t,t+1}$ is defined as
$E_{t,t+1}=\mathcal{V}_{t}^{\prime}\times\mathcal{V}_{t+1}^{\prime}.$ (7)
The process progresses iteratively until the stop condition is met, and finally, we get a T-FFN $\mathcal{G}_{q}^{\mathcal{A}}$ for the query input $q$. We use function $f_{\mathrm{IAS}}$ to denote the entire computation:
$\mathcal{G}_{q}^{\mathcal{A}}=f_{\mathrm{IAS}}(\mathcal{A},q).$ (8)